if P(A) =0.45, then P(A) = ?
1
Expert's answer
2019-08-22T11:49:12-0400

Given that "P(A)=0.45, P(B)=0.35, P(A\cap B)=0.25"

Complement rule

"P(\bar{A})=1-P(A)=1-0.45=0.55"

Addition rule


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"P(A\cup B)=0.45+0.35-0.25=0.55"